OpenStudy (timaashorty):

HELP PLEASE Use the quadratic formula to solve the equation. x^2-7x-6=0

OpenStudy (psymon):

You know the formula and what to plug in?

OpenStudy (timaashorty):

ax2 + bx + c = 0 is the formula ?

OpenStudy (psymon):

\[\frac{ -b \pm \sqrt{b^{2} - 4ac} }{ 2a }\]

OpenStudy (anonymous):

psymon's right

OpenStudy (anonymous):

a would be 1, b is -7 then c is -6

OpenStudy (timaashorty):

Okay thats the formula, how would I solve it ?

OpenStudy (anonymous):

substitute it in the formula..

OpenStudy (timaashorty):

Can you show me how ?

OpenStudy (anonymous):

\[\frac{ 7\pm \sqrt{(-7)^2 - 4(1)(-6)} }{ 2(1) }\]

OpenStudy (anonymous):

there would be 2 answers..

OpenStudy (anonymous):

\[7+\frac{\sqrt{73} }{ 2 }\]

OpenStudy (anonymous):

and \[7- \frac{ \sqrt{73}}{ 2 }\]

OpenStudy (anonymous):

thats it..

OpenStudy (timaashorty):

omg, thank you so much silver that helped me alot!

OpenStudy (anonymous):

youre welcome :)